OpenStudy (anonymous):

PLZ HELP WILL MEDAL!!!!!!!!!! A line has a slope of -1/2 and a y-intercept of 3. Which equations represent the line? Choose all answers that are correct.

OpenStudy (anonymous):

A. 2Y=6+X B. X+2Y=6 C. Y=-1/2X+3 D. (Y-2)=-1/2(X-2)

OpenStudy (anonymous):

slope is -1/2 any y intercept is 3 so its suppose to come out looking like y=slopex+y so for example slope -2 and y intercept 4 would look like y=-2x+4

OpenStudy (anonymous):

@lollypop12345 might as well give the definition of the slope-intercept formula which is: \[y = mx + c\] Where: \(m =\) slope. \(c =\) y-intercept. \(c\) may be a different letter depending on which country you are from.

OpenStudy (anonymous):

thats where i was trying to get at thanks but i cant say the answer its my first day on open study i know how to do it because i took college algebra in college and we learnt it

OpenStudy (anonymous):

No worries mate, you did well. Keep up the good work.